OpenStudy (anonymous):

Which value of x makes the open sentence true? 7 – x = 3x + 5 A. x = 2 B. x = 0.5 C. x = 3 D. x = 1.5

OpenStudy (anonymous):

i think it might be A

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

yeahhh

OpenStudy (anonymous):

7 - 5 = 3x + x 2 = 4x 1/2 = x 0.5 = x

OpenStudy (anonymous):

so now we have a debate which one is right

OpenStudy (anonymous):

dang i was wrong

OpenStudy (anonymous):

hey this one guy needs my help brb